OpenStudy (anonymous):

If f(x) = x - 4, what does f(3) represent?

OpenStudy (anonymous):

The value of (x - 4) when x = 3 The value if x when (x - 4) = 3 The value of 3f(x) when x = 3 The value of f(x) when (x - 4) = 3

hartnn (hartnn):

to get f(3) from f(x), just plug in x= 3 in f(x)!

hartnn (hartnn):

plugging in x=3 means you're finding the value of x-4 when x= 3 :P

OpenStudy (anonymous):

A.?

hartnn (hartnn):

yes!

OpenStudy (anonymous):

Thank you!

hartnn (hartnn):

welcome ^_^